OpenStudy (anonymous):

1) Find the domain and range for the relation: 1. {(2,7),(4,9),(2,11),(7,13),(5,9)}

OpenStudy (anonymous):

oh i was just about two say that , i have a few me question

OpenStudy (anonymous):

can you help me?

OpenStudy (anonymous):

some is and the rest are function

OpenStudy (anonymous):

here one: 3) Evaluate the function rule for x = -4: 1. F(x) = 15 - x

OpenStudy (anonymous):

ok thanks for taking the time out and helping me... 4) Find the range of function rule y = 3x + 4 for the given domain: {0,5,7,11}

OpenStudy (anonymous):

i have 6 more questions after he one that i already posted

OpenStudy (aripotta):

if the domain is all the x's, then just substitute for x with each of the givens

OpenStudy (aripotta):

for 0: y = 3x + 4. y = 3(0) + 4. y = 4

OpenStudy (aripotta):

can you do the rest?

OpenStudy (anonymous):

yea thanks ok i have some more

OpenStudy (anonymous):

HERE: 5) Write a function rule for each situation: 1. The total cost of pounds of pears if each pound costs $1.19 2. The amount of money you can make shoveling snow if you receive $8.00 per driveway

OpenStudy (aripotta):

y = 1.19x y = 8x

OpenStudy (anonymous):

OK THANKS NEXT ONE 6) Find the common difference of the arithmetic sequence: -2, -6, -10, -14…

OpenStudy (aripotta):

what do you think?

OpenStudy (anonymous):

I DON'T NO I REALLY DON'T KNOW

OpenStudy (aripotta):

what's the difference between -2 and -6?

OpenStudy (anonymous):

um -2 is is bigger

OpenStudy (aripotta):

no i mean, -2 minus what equals -6?

OpenStudy (anonymous):

4

OpenStudy (aripotta):

so that's the answer :)

OpenStudy (anonymous):

so the answeer is 4

OpenStudy (aripotta):

yes, it's the common difference

OpenStudy (anonymous):

ok another one.

OpenStudy (anonymous):

here: 7) Find the third term of each arithmetic sequence: A(n) = 8 + (n-1)(-3)

OpenStudy (aripotta):

don't you just substitute 3 for n?

OpenStudy (anonymous):

im not sure

OpenStudy (aripotta):

i think that's what you do

OpenStudy (anonymous):

yea i just look it up so can you help

OpenStudy (aripotta):

8 + (3-1)(-3)

OpenStudy (anonymous):

can you help me work it right

OpenStudy (anonymous):

out

OpenStudy (aripotta):

3 - 1 = ?

OpenStudy (anonymous):

4

OpenStudy (aripotta):

it would be 2 lol

OpenStudy (aripotta):

2(-3) = ?

OpenStudy (anonymous):

-6

OpenStudy (aripotta):

8 - 6 = ?

OpenStudy (anonymous):

2

OpenStudy (aripotta):

so 2 is your answer

OpenStudy (anonymous):

2?

OpenStudy (aripotta):

yea

OpenStudy (anonymous):

hold on one sec can yu help someone else for a minute so i can right this down

OpenStudy (aripotta):

ok

OpenStudy (anonymous):

2. A(n) = -3(n-2) + 4

OpenStudy (aripotta):

we're trying to find the third term again, right?

OpenStudy (aripotta):

-3(3-2) + 4

OpenStudy (anonymous):

ok so then it would be -3(1)+4 is that right

OpenStudy (aripotta):

yes

OpenStudy (anonymous):

so it would be -3+4 so the answer is 1

OpenStudy (aripotta):

yes

OpenStudy (anonymous):

ok give me a little time again right back ok

OpenStudy (anonymous):

8) Use the sequence -4, 8, -16, 32… 1. What is the common ratio? 2. What are the next three terms?

OpenStudy (aripotta):

i don't know what it means by ratio, but it's probably -2

OpenStudy (anonymous):

ok what about the next three terms would it be -64, -128, -256

OpenStudy (aripotta):

-128 should be positive

OpenStudy (anonymous):

ok so i have one more

OpenStudy (aripotta):

mmk

OpenStudy (anonymous):

here: 9) Critical Thinking: Buses on your route run every 9 minutes from 6:00 AM to 10:00 AM. You get to the bus stop at 7:16 AM,… a. What is the maximum amount of time you should have to wait for the bus? b. How many minutes after the buses begin running at 6:00 AM, do you arrive at the bus stop? c. What time does the first bus of the day arrive at the bus stop?

OpenStudy (aripotta):

uh

OpenStudy (anonymous):

9) Critical Thinking: Buses on your route run every 9 minutes from 6:00 AM to 10:00 AM. You get to the bus stop at 7:16 AM,… a. What is the maximum amount of time you should have to wait for the bus? b. How many minutes after the buses begin running at 6:00 AM, do you arrive at the bus stop? c. What time does the first bus of the day arrive at the bus stop?

OpenStudy (anonymous):

you don't get it

OpenStudy (aripotta):

not really

OpenStudy (anonymous):

ok well do you know anyone that would no this and heres you medal

OpenStudy (aripotta):

@Hero maybe?

OpenStudy (anonymous):

ok thanks talk to you lter and if you want you can give me a medal for asking the question

OpenStudy (anonymous):

thanks bye!

OpenStudy (aripotta):

bai and gl